# Reference-based Metrics Disprove Themselves in Question Generation

### v2

@@ Abstract
Reference-based metrics such as BLEU and BERTScore are widely used to evaluate question generation (QG). In this study, on QG benchmarks such as SQuAD and HotpotQA, we find that using human-written references cannot guarantee the effectiveness of the reference-based metrics. Most QG benchmarks have only one reference; we replicated the annotation process and collect another reference. A good metric was expected to grade a human-validated question no worse than generated questions. However, the results of reference-based metrics on our newly collected reference disproved the metrics themselves. We propose a reference-free metric consisted of multi-dimensional criteria such as naturalness, answerability, and complexity, utilizing large language models. These criteria are not constrained to the syntactic or semantic of a single reference question, and the metric does not require a diverse set of references. Experiments reveal that our metric accurately distinguishes between high-quality questions and flawed ones, and achieves state-of-the-art alignment with human judgment.
@@ 1 Introduction
To overcome the limitation of the answerability measure in RQUGE and implement the other two measures, we utilize large language models (LLMs), which have demonstrated potential utility in data annotation tasks Liu et al. (2023); Wang et al. (2023); Lin and Chen (2023); Chiang and Lee (2023), and their Chain-of-Thought (CoT) Wei et al. (2022) process. We design CoT prompts for the LLM to directly measure the three criteria, as described in detail in §$\S$3.
@@ 2.1 Study Design & Data Collection
Previous studies questionting the effectiveness of reference-based metrics in QG typically rely on human evaluation. That is, they investigate whether the scores given to generated questions by QG metrics are highly correlated with the scores given by human evaluators Mohammadshahi et al. (2023); Ji et al. (2022). Unlike these studies, our research adopts a different approach during the data collection phase for QG datasets. Specifically, we replicate the data collection procedure of the datasets to collect new references, referred to as Group 1. Our focus is on determining if the newly collected references, when evaluated as candidates against the original references, receive high ratings from existing metrics. In addition, we extended our collection procedure to include three additional groups of candidate questions considered less desirable (Groups 2, 3, and 4) to ensure comprehensive comparisons. An effective and robust evaluation metric should assign a significantly higher score for questions in Group 1 compared to those in other groups. Fig. 1 illustrates our data collection process.
@@
Liu et al. proposed a concept of clues for QG, which refers to words from the context passage that also appear in the question. Their experimental results indicate that the addition of a clue-prediction model enhances the performance of question generators on reference-based metrics. We investigate the usefulness of this concept by asking the annotators to phrase an additional question such that it contains the clue words used by the original annotators of the datasets. We ensure that the clue words are only presented to the annotators after they have finished asking their first question.
We perform the additional annotation on two popular QG benchmarks: (1) 748 test examples of SQuAD Rajpurkar et al. (2016), and (2) 48 test examples of HotpotQA Yang et al. (2018). To illustrate the application of our study, we collect another QG dataset in the educational domain, specifically from the TED-Ed learning platform [^1] https://ed.ted.com/. We further annotate 43 questions from this new dataset. More details about data collection and annotation for TedEdQA are provided in Appx. A.5.
For the HotpotQA sample, we also collected three other sets of questions, each violating an aspect required by the reference questions.
Group 2: Single-hop questions for a multi-hop QG benchmark: This group of candidate questions targets the multi-hop characteristic of HotpotQA where the ground-truth questions are formed based on two passages. Specifically, we selected one from the original two passages that contains the answer span. We then asked GPT-3.5 to generate a question based on this single passage. We reviewed the questions for grammar, clarity, relevance to the passage, independence from external knowledge, and a logical path to the answer.
Group 3: Non-questions that use the same words as the reference: For this group of questions, we asked GPT-3.5 to generate a sentence based on the passages and use as many words from the same list of clues given to our annotators. We added a constraint such that the generated sentence cannot be in the form of a question. We then manually went through the generated sentences to ensure that no hallucinations were in place. In this sense, we produced a group of candidates that does not satisfy the most basic linguistic requirement of a question, naturalness, but still manages to contain many similar words as the ground-truth questions.
@@ 2.2 Results
Fig. 1 shows the average normalized scores given by reference-based metrics to the four groups of candidate questions, all based on the same references. We find that all reference-based metrics, BLEU, ROUGE-L, BLEURT, Q-BLEU, and BERTScore, can effectively distinguish Group 4 (random questions) from the other groups, assigning it significantly lower score. For instance, the average ROUGE-L score for Group 4 is $0.11$, compared to $0.31$ for Group 1, $0.25$ for Group 2, and $0.39$ for Group 3, with a minimum difference of $14\%$ from the scores of the other groups.
Fig. 1 also reveals issues with the reference-based metrics in accurately assessing Groups 1, 2, and 3. Notably, for all five reference-based metrics, Group 3, non-question sentences with wording similar to the references, receives the highest average score. For example, the ROUGE-L metric scores a non-question sentence that uses similar wording to the reference (green bar) on average $8\%$ higher than a new reference produced by our annotators (red bar), and $14\%$ higher than a perfectly answerable question requiring less reasoning than the reference (blue bar). This observation indicates a flaw in reference-based metrics, as candidates that do not form coherent questions should not receive higher scores than those that do.
The recently-introduced reference-free metrics, QAScore and RQUGE, also face difficulties in giving reasonable scores to questions from Groups 1, 2 and 3. QAScore, despite rating the new references highest among four groups, shows minimal score differences. Meanwhile, RQUGE gives the highest average score ($0.85$) to Group 2, which contains single-hop questions in contexts requiring multi-hop reasoning. RQUGE’s preference for single-hop questions can be attributed to its disregard for the complexity of the candidate question. It utilizes a pretrained QA model to compute a score based on the model’s responses to the candidate question. The questions we collected, which require reasoning over two documents, may pose a greater challenge for the QA model compared to the simpler questions from Group 2. Since RQUGE’s scoring mechanism does not consider the question’s complexity, it underestimates the new references we collected in Group 1, scoring them at $0.80$.
@@ 3 NACo: A Novel Multi-dimensional Reference-free QG Metric
  • The LLM first reads over the context passage(s) and the question. The LLM check whether the question makes these mistakes: (1) not a question, (2) grammar errors, or (3) unclear objective. If so, the LLM should respond with ‘Question unnatural’, and we assign a score of 00 for the question in terms of naturalness $n_{cand}$. Otherwise, $n_{cand}$ is $1$.
@@
  • The LLM provide the final answer to the question. We define the answerability of the question $a_{cand}$ as the F1 score between the LLM’s answer to the question and the ground-truth answer used to generate the question.
@@
If a candidate question passes the initial naturalness and answerability evaluation, we determine whether its complexity aligns with expected standards for the domain and dataset. For example, in the HotpotQA dataset, questions that require multi-hop reasoning might be preferred over simpler, single-hop questions. This preference may not hold in other datasets. In this sense, NACo relies on a subset of examples from the specific dataset to find the expected complexity of a question in that dataset. Specifically, we perform the above CoT-QA process to obtain the complexity of the references. Expected complexity is then defined by the most common number of reasoning steps needed by the LLM to answer a reference question. Our experiments using 750 examples from the training set find that for GPT-3.5, the expected complexity is $2$ for questions in SQuAD, and $3$ for questions in HotpotQA. Subsequently, NACo measures the similarity, denoted by $c_{cand}$, between the complexity of the candidate question and the expected complexity.
@@ B B-RT R-L BSc Q-B
       B -0.26 -0.24 -0.20 -0.19 -0.17 -0.15 0.00 0.02 0.02 -0.13 -0.07 -0.06
         w/ DivRef 0.09 0.08 0.07 0.14 0.13 0.11 0.29 0.28 0.24 0.27 0.26 0.22
       B-RT 0.11 0.09 0.07 0.16 0.14 0.11 0.33 0.33 0.25 0.30 0.28 0.21
         w/ DivRef 0.14 0.12 0.10 0.20 0.20 0.16 0.36 0.37 0.29 0.34 0.34 0.25
       R-L 0.09 0.04 0.03 0.11 0.06 0.05 0.30 0.27 0.22 0.25 0.19 0.15
         w/ DivRef 0.16 0.17 0.13 0.20 0.19 0.16 0.36 0.36 0.29 0.34 0.33 0.25
       BSc 0.19 0.18 0.15 0.21 0.22 0.18 0.39 0.41 0.33 0.37 0.38 0.28
         w/ DivRef 0.24 0.25 0.19 0.29 0.31 0.24 0.44 0.46 0.37 0.44 0.46 0.34
       Q-B 0.09 0.05 0.04 0.13 0.08 0.07 0.33 0.34 0.26 0.28 0.25 0.18
         w/ DivRef 0.10 0.05 0.04 0.14 0.11 0.09 0.33 0.34 0.26 0.27 0.26 0.19
@@
       Llama3-8B 0.43 0.33 0.27 0.60 0.51 0.43 0.44 0.36 0.29 0.65 0.49 0.38
       Mixtral-8x7B 0.37 0.29 0.23 0.58 0.52 0.43 0.41 0.27 0.21 0.59 0.43 0.33
       Claude-Haiku 0.47 0.26 0.21 0.67 0.49 0.40 0.53 0.44 0.36 0.73 0.54 0.43
       GPT3.5 0.40 0.25 0.19 0.70 0.59 0.49 0.44 0.34 0.27 0.68 0.49 0.38
       GPT4 0.55 0.29 0.25 0.72 0.56 0.48 0.56 0.38 0.36 0.8 0.53 0.44
@@ 4.1 Experimental Setup
Baselines: We compare the evaluation capacity of NACo with five reference-based metrics, including BLEU-4 (B) Papineni et al. (2002), BLEURT (B-RT) Sellam et al. (2020), ROUGE-L (R-L) Lin (2004), BERTScore (BSc) Zhang et al. (2019), and Q-BLEU (Q-B) Nema and Khapra (2018), and two reference-free metrics, QAScore (QA-S) Ji et al. (2022), and RQUGE (R-Q)Mohammadshahi et al. (2023). Tang et al. propose using LLM to diversify the limited references in benchmarks, demonstrating an improvement in the correlation between reference-based metrics and human judgment. We replicate this approach and report the evaluation performance of the five reference-based metrics both when only the original reference is used and when adding the diversified references.
@@
Human Evaluation: We recruit volunteer annotators, all fluent English speakers, to evaluate both model-generated questions and human-written questions, using 48 test examples from HotpotQA. For each example, annotators evaluate four questions: RefQ, GPT-3.5 (zero-shot), BART-base, and AnnoQ, displayed in randomized and anonymized order. Evaluators rate each question based on naturalness, answerability, and complexity, using a 3-point scale for each criterion. Additionally, we sum the individual scores together to calculate a combined score that reflects the question’s overall quality. We obtain three annotations per question and use the average of these as the standard for human judgment. The Pearson correlation coefficient between ratings given by our annotators are 0.67. The rating rubric is available in Appx. A.4.1.
@@ 4.2 Results
Failure of reference-based metrics: We report QG competitors’ performance on various metrics, including NACo, using RefQ as the reference in Tbl. 1 for SQuAD. Even though RefQ, AnnoQ, and AnnoQ-clue-RefQ are all qualified as valid questions, reference-based metrics rate them with significant differences. In the SQuAD dataset, BLEU scores for RefQ, AnnoQ, and AnnoQ-clue-RefQ are 100, 12.78, and 27.43, respectively (Tbl. 1). However, NACo score these three groups questions similarly, with RefQ, AnnoQ, and AnnoQ-clue-RefQ scoring 75.09, 74.01, and 74.21, respectively (Tbl. 1). Similar patterns are observed in the HotpotQA and TedEdQA datasets, as detailed in Tbl. 6 and Tbl. 7.
@@
Effectiveness of NACo: Referring to our analysis of four groups of candidate questions for HotpotQA in Fig. 1, NACo uniquely succeeds in separating all four groups by significant margins, unlike the seven existing metrics. The newly collected multi-hop questions in Group 1, which satisfy all criteria for HotpotQA questions, achieve the highest average NACo score of $0.83$. They are followed by the questions in Group 2, lacking in complexity, with a score of $0.75$; Group 3, lacking in naturalness, with a score of $0.19$; and Group 4, lacking in answerability, with a score of $0.03$.
We calculate the Pearson $r$, Spearman $\rho$, and Kendall $\tau$ correlation coefficients to measure the agreement between all metrics, including NACo, and human judgment, as reported in Tbl. 2. This comparison considers correlation with both individual criteria and the overall question quality. Tbl. 2 reveals that NACo demonstrates the highest correlation with human evaluation for individual criteria in 7 out of 9 scores, and it ranks second-highest in the remaining 2 scores. Notably, NACo exhibits the strongest agreement with human judgment concerning the overall quality of questions across all correlation metrics. This observation is consistent across different underlying LLMs.

### v3

@@ Abstract
Reference-based metrics such as BLEU and BERTScore are widely used to evaluate question generation (QG). In this study, on QG benchmarks such as SQuAD and HotpotQA, we find that using human-written references cannot guarantee the effectiveness of the reference-based metrics. Most QG benchmarks have only one reference; we replicate the annotation process and collect another reference. A good metric is expected to grade a human-validated question no worse than generated questions. However, the results of reference-based metrics on our newly collected reference disproved the metrics themselves. We propose a reference-free metric consisted of multi-dimensional criteria such as naturalness, answerability, and complexity, utilizing large language models. These criteria are not constrained to the syntactic or semantic of a single reference question, and the metric does not require a diverse set of references. Experiments reveal that our metric accurately distinguishes between high-quality questions and flawed ones, and achieves state-of-the-art alignment with human judgment.
@@ 1 Introduction
To overcome the limitation of the answerability measure in RQUGE (Mohammadshahi et al., 2023) and implement the other two measures, we utilize large language models (LLMs), which have demonstrated potential utility in data annotation tasks Liu et al. (2023); Wang et al. (2023); Lin and Chen (2023); Chiang and Lee (2023), and their Chain-of-Thought (CoT) Wei et al. (2022) process. We design CoT prompts for the LLM to directly measure the three criteria, as described in detail in §$\S$3.
@@
We release the collected data and code implementation of NACo to facilitate future works.  [^1] https://github.com/bnguyen5/naco
@@ 2.1 Study Design & Data Collection
Previous studies questioning the effectiveness of reference-based metrics in QG typically rely on human evaluation. That is, they investigate whether the scores given to generated questions by QG metrics are highly correlated with the scores given by human evaluators Mohammadshahi et al. (2023); Ji et al. (2022). Unlike these studies, our research adopts a different approach during the data collection phase for QG datasets. Specifically, we replicate the data collection procedure of the datasets to collect new references, referred to as Group 1. Our focus is on determining if the newly collected references, when evaluated as candidates against the original references, receive high ratings from existing metrics. In addition, we extended our collection procedure to include three additional groups of candidate questions considered less desirable (Groups 2, 3, and 4) to ensure comprehensive comparisons. An effective and robust evaluation metric should assign a significantly higher score for questions in Group 1 compared to those in other groups. Fig. 1 illustrates our data collection process.
@@
Liu et al. proposed a concept of clues for QG, which refers to words from the context passage that also appear in the question. Their experimental results indicate that the addition of a clue-prediction model enhances the performance of question generators on reference-based metrics. We investigate the usefulness of this concept by asking the annotators to phrase an additional question such that it contains the clue words used by the original annotators of the datasets. We ensure that the clue words are only presented to the annotators after they have finished creating their first question.
We perform the additional annotation on two popular QG benchmarks: (1) 748 test examples of SQuAD Rajpurkar et al. (2016), and (2) 96 test examples of HotpotQA Yang et al. (2018). To illustrate the application of our study, we collect another QG dataset in the educational domain, specifically from the TED-Ed learning platform [^2] https://ed.ted.com/. We further annotate 43 questions from this new dataset. More details about data collection and annotation for TedEdQA are provided in Appx. A.5.
For the HotpotQA sample, we also collect three other sets of questions, each violating an aspect required by the reference questions.
Group 2: Single-hop questions for a multi-hop QG benchmark: This group of candidate questions targets the multi-hop characteristic of HotpotQA where the ground-truth questions are formed based on two passages. Specifically, we select one from the original two passages that contains the answer span. We then ask GPT-3.5 to generate a question based on this single passage. We review the questions for grammar, clarity, relevance to the passage, independence from external knowledge, and a logical path to the answer.
Group 3: Non-questions that use the same words as the reference: For this group of questions, we ask GPT-3.5 to generate a sentence based on the passages and use as many words from the same list of clues given to our annotators. We add a constraint such that the generated sentence cannot be in the form of a question. We then manually go through the generated sentences to ensure that no hallucinations were in place. In this sense, we produce a group of candidates that does not satisfy the most basic linguistic requirement of a question, naturalness, but still manages to contain many similar words as the ground-truth questions.
@@ 2.2 Results
Fig. 1 shows the average normalized scores given by reference-based metrics to the four groups of candidate questions, all based on the same references. We find that all reference-based metrics, BLEU, ROUGE-L, BLEURT, Q-BLEU, and BERTScore, can effectively distinguish Group 4 (random questions) from the other groups, assigning it significantly lower scores. For instance, the average ROUGE-L score for Group 4 is $0.10$, compared to $0.34$ for Group 1, $0.26$ for Group 2, and $0.40$ for Group 3, with a minimum difference of $16\%$ from the scores of the other groups.
Fig. 1 also reveals issues with the reference-based metrics in accurately assessing Groups 1, 2, and 3. Notably, for all five reference-based metrics, Group 3, non-question sentences with wording similar to the references, receives the highest average score. For example, the ROUGE-L metric scores a non-question sentence that uses similar wording to the reference (green bar) on average $6\%$ higher than a new reference produced by our annotators (red bar), and $14\%$ higher than a perfectly answerable question requiring less reasoning than the reference (blue bar). This observation indicates a flaw in reference-based metrics, as candidates that do not form coherent questions should not receive higher scores than those that do.
The recently-introduced reference-free metrics, QAScore and RQUGE, also face difficulties in giving reasonable scores to questions from Groups 1, 2 and 3. QAScore, despite rating the new references highest among four groups, shows minimal score differences. Meanwhile, RQUGE gives the highest average score ($0.84$) to Group 2, which contains single-hop questions in contexts requiring multi-hop reasoning. RQUGE’s preference for single-hop questions can be attributed to its disregard for the complexity of the candidate question. It utilizes a pretrained QA model to compute a score based on the model’s responses to the candidate question. The questions we collected, which require reasoning over two documents, may pose a greater challenge for the QA model compared to the simpler questions from Group 2. Since RQUGE’s scoring mechanism does not consider the question’s complexity, it underestimates the new references we collected in Group 1, scoring them at $0.77$.
@@ 3 NACo: A Novel Multi-dimensional Reference-free QG Metric
  • The LLM first reads over the context passage(s) and the question. The LLM checks whether the question makes these mistakes: (1) not a question, (2) grammar errors, or (3) unclear objective. If so, the LLM should respond with ‘Question unnatural’, and we assign a score of 00 for the question in terms of naturalness $n_{cand}$. Otherwise, $n_{cand}$ is $1$.
@@
  • The LLM provides the final answer to the question. We define the answerability of the question $a_{cand}$ as the F1 score between the LLM’s answer to the question and the ground-truth answer used to generate the question.
@@
If a candidate question passes the initial naturalness and answerability evaluation, we determine whether its complexity aligns with expected standards for the domain and dataset. For example, in the HotpotQA dataset, questions that require multi-hop reasoning might be preferred over simpler, single-hop questions. This preference may not hold in other datasets. In this sense, NACo relies on a subset of examples from the specific dataset to find the expected complexity of a question in that dataset. Specifically, we perform the above CoT-QA process to obtain the complexity of the references. Expected complexity is then defined by the most common number of reasoning steps needed by the LLM to answer a reference question. In our experiments, we use 750 examples from the training set of SQuAD and HotpotQA to compute the expected complexity for each dataset. Subsequently, NACo measures the similarity, denoted by $c_{cand}$, between the complexity of the candidate question and the expected complexity.
@@ B B-RT R-L BSc Q-B
       B 0.08 -0.06 -0.04 0.16 0.06 0.05 0.40 0.38 0.31 0.25 0.20 0.16
         w/ DivRef 0.09 -0.01 -0.01 0.18 0.10 0.08 0.43 0.42 0.33 0.27 0.25 0.19
       B-RT 0.12 0.03 0.03 0.23 0.19 0.14 0.49 0.50 0.37 0.33 0.35 0.24
         w/ DivRef 0.14 0.05 0.04 0.26 0.22 0.16 0.51 0.52 0.39 0.36 0.38 0.26
       R-L 0.07 -0.03 -0.03 0.18 0.10 0.07 0.43 0.42 0.31 0.27 0.24 0.17
         w/ DivRef 0.13 0.06 0.04 0.24 0.19 0.15 0.48 0.48 0.37 0.34 0.34 0.24
       BSc 0.18 0.09 0.06 0.27 0.21 0.16 0.52 0.54 0.41 0.38 0.39 0.27
         w/ DivRef 0.23 0.14 0.10 0.33 0.29 0.21 0.56 0.57 0.43 0.44 0.45 0.31
       Q-B 0.07 -0.05 -0.04 0.17 0.09 0.07 0.45 0.44 0.32 0.27 0.26 0.17
         w/ DivRef 0.07 -0.04 -0.03 0.16 0.09 0.07 0.45 0.44 0.32 0.27 0.26 0.18
@@
       Llama3-8B 0.49 0.32 0.25 0.66 0.50 0.39 0.43 0.36 0.21 0.64 0.42 0.30
       Mixtral-8x7B 0.36 0.30 0.23 0.54 0.52 0.40 0.36 0.23 0.17 0.56 0.40 0.28
       Claude-Haiku 0.53 0.30 0.23 0.71 0.51 0.40 0.50 0.35 0.27 0.71 0.47 0.35
       GPT3.5 0.47 0.30 0.23 0.70 0.53 0.41 0.49 0.35 0.27 0.68 0.48 0.35
       GPT4 0.64 0.36 0.28 0.78 0.59 0.47 0.55 0.35 0.27 0.80 0.52 0.39
@@ 4.1 Experimental Setup
Baselines: We compare the evaluation capacity of NACo with five reference-based metrics, including BLEU-4 (B) Papineni et al. (2002), BLEURT (B-RT) Sellam et al. (2020), ROUGE-L (R-L) Lin (2004), BERTScore (BSc) Zhang et al. (2019), and Q-BLEU (Q-B) Nema and Khapra (2018), and two reference-free metrics, QAScore (QA-S) Ji et al. (2022), and RQUGE (R-Q)Mohammadshahi et al. (2023). Tang et al. proposes using LLM to diversify the limited references in benchmarks, demonstrating an improvement in the correlation between reference-based metrics and human judgment. We replicate this approach and report the evaluation performance of the five reference-based metrics both when only the original reference is used and when adding the diversified references.
@@
Human Evaluation: We recruit volunteer annotators, all fluent English speakers, to evaluate both model-generated questions and human-written questions, using 96 test examples from HotpotQA. For each example, annotators evaluate four questions: RefQ, GPT-3.5 (zero-shot), BART-base, and AnnoQ, displayed in randomized and anonymized order. Evaluators rate each question based on naturalness, answerability, and complexity, using a 3-point scale for each criterion. Additionally, we sum the individual scores to calculate a combined score that reflects the question’s overall quality. We obtain three annotations per question and use the average of these as the standard for human judgment. The Pearson correlation coefficient between ratings given by our annotators is 0.67. The rating rubric is available in Appx. A.4.1.
@@ 4.2 Results
Failure of reference-based metrics: We report QG competitors’ performance on various metrics, including NACo, using RefQ as the reference in Tbl. 1 for SQuAD. Even though RefQ, AnnoQ, and AnnoQ-clue-RefQ are all qualified as valid questions, reference-based metrics rate them with significant differences. In the SQuAD dataset, BLEU scores for RefQ, AnnoQ, and AnnoQ-clue-RefQ are 100, 12.78, and 27.43, respectively (Tbl. 1). However, NACo rates these three groups of questions similarly, with RefQ, AnnoQ, and AnnoQ-clue-RefQ scoring 75.09, 74.01, and 74.21, respectively (Tbl. 1). Similar patterns are observed in the HotpotQA and TedEdQA datasets, as detailed in Tbl. 7 and Tbl. 8.
@@
Effectiveness of NACo: Referring to our analysis of four groups of candidate questions for HotpotQA in Fig. 1, NACo uniquely succeeds in separating all four groups by significant margins, unlike the seven existing metrics. The newly collected multi-hop questions in Group 1, which satisfy all criteria for HotpotQA questions, achieve the highest average NACo score of $0.85$. They are followed by the questions in Group 2, lacking in complexity, with a score of $0.80$; Group 3, lacking in naturalness, with a score of $0.21$; and Group 4, lacking in answerability, with a score of $0.02$.
We calculate the Pearson $r$, Spearman $\rho$, and Kendall $\tau$ correlation coefficients to measure the agreement between all metrics, including NACo, and human judgment, as reported in Tbl. 2. This comparison considers correlation with both individual criteria and the overall question quality. Tbl. 2 reveals that NACo demonstrates the highest correlation with human evaluation for individual criteria in 9 out of 12 scores. Notably, NACo exhibits the strongest agreement with human judgment concerning the overall quality of questions across all correlation metrics. This observation is consistent across different underlying LLMs.
4.3 Analysis